OpenStudy (anonymous):

A ground-ground projectile launched from O with a speed u has to clear an obstacle PQ on its path which is 30m height and a horizontal distance of 40 m . Find the min u required

OpenStudy (anonymous):



OpenStudy (anonymous):

tanx = 30/40 tanx=3/4 is that correct!

OpenStudy (experimentx):